# Fact-Checking with Contextual Narratives: Leveraging Retrieval-Augmented LLMs for Social Media Analysis

### v1

@@ Abstract.
Machine Learning, Fact-Checking, LLMs, Retrieval, Clustering
ccs: Information systems Expert systems
Figure 1. CRAVE is a fact-checking method that analyses multimodal claims, expressed in social media posts, to determine their veracity. It retrieves multimodal evidence, clusters it into distinct narratives, and uses an LLM with 5W1H (who, what, when, where, why, and how) reasoning to produce an interpretable explanation and veracity verdict.
1. Introduction
Fact-checking has emerged as a critical task in the era of social media, where information spreads rapidly(Vosoughi et al., 2018; Aïmeur et al., 2023; Cooke, 2017) and misinformation can have far-reaching consequences(Kshetri and Voas, 2017; Caceres et al., 2022; Allcott and Gentzkow, 2017). Automated fact-checking systems have gained traction as scalable solutions, yet they often grapple with challenges such as handling diverse evidence sources, integrating multimodal data, and presenting comprehensive narratives. Traditional approaches (Papadopoulos et al., 2024a, b) focus on verifying factual claims against a static repository of information(Papadopoulos et al., 2024c; Luo et al., 2021; Abdelnabi et al., 2022; Tonglet et al., 2025), which may overlook nuances in evidence and context.
Recent advances in Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) have opened new avenues for building robust fact-checking frameworks. RAG has been shown effective in reducing factual inconsistencies (”hallucinations”) during text generation, making it attractive for fact-checking task(Lewis et al., 2020; Momii et al., 2024). These models can access dynamic external knowledge, reason over complex evidence, and produce context-aware summaries. However, effectively leveraging these capabilities requires addressing several challenges: (1) how to retrieve relevant evidence efficiently, (2) how to cluster and organize potentially contradictory bodies of evidence into coherent narratives, and (3) how to generate holistic evaluations that incorporate these multiple perspectives.
In this work, we propose CRAVE (Cluster-based Retrieval Augmented Verification with Explanation); a novel framework that integrates retrieval-augmented LLMs with clustering techniques to identify, structure, and evaluate evidence for fact-checking social media posts. The framework is designed to process multi-modal inputs (text and images) and iteratively refine evidence through agent-based mechanisms. By clustering evidence into distinct narratives, and identifying external support (or otherwise) for those narratives, CRAVE enables the assessment of contentious topics that present with support from contradictory and noisy news sources. The contributions of CRAVE are threefold:
@@ 2. Related work
Automated fact-checking systems deployed in practice commonly work by matching newly surfaced claims against a library of already fact-checked claims (Graves, 2018). While efficient for well-known misinformation that periodically resurfaces, such systems struggle with novel or emerging misinformation. Research to date has mostly focused upon detecting intrinsic artifacts within a claim (Liu et al., 2024b; Shao et al., 2023) using synthetically generated datasets (Aneja et al., 2021; Abdelnabi et al., 2022; Shao et al., 2023; Liu et al., 2024b). Such black box detection methods often lack explainability and lag behind development in generative AI, so limiting their effectiveness and on novel misinformation.
@@
Evidence-driven methods and datasets have been introduced to enhance the explainability of fact-checking frameworks. However, the evidence used is often curated and sanitized (Papadopoulos et al., 2024c; Abdelnabi et al., 2022). As a result, methods such as (Papadopoulos et al., 2024b, a) achieve state-of-the-art results while bypassing the challenge of retrieving relevant evidence, which limits their applicability to real-world scenarios. Although these approaches are evidence-based, they are primarily data-driven trained classifiers that struggle to generalize across novel claims.
2.1. Retrieval of evidence
@@
Effective verification of multimodal misinformation depends on locating reliable evidence, particularly for images reused out of context (Abdelnabi et al., 2022; Papadopoulos et al., 2024a; Braun et al., 2025). As identified in (Graves, 2018) evidence retrieval is still done primarily by human fact-checkers through manually constructing search terms. Prior works have demonstrated that incorporating reverse image search to find an image’s provenance and text-based search for contextual information can similarly improve automated fact-checking performance (Abdelnabi et al., 2022; Tonglet et al., 2025; Dey et al., 2024; Qi et al., 2024). Abdelnabi et al. were among the first to show that combining reverse image search with text retrieval improves scrutiny of multimedia claims. Since then, other pipelines have extended such concepts to incorporate more robust retrieval modules or specialized knowledge. In (Tonglet et al., 2025) the authors extend their sources through WIKIDATA(Vrandečić and Krötzsch, 2014) and Wikichat(Semnani et al., 2023) to enrich their image context, but the authors also note the limited effectiveness in non-Western contexts.
@@
In real-world scenarios, fact-checking is complicated by how social media users often combine verbose or biased commentary with images unrelated to mainstream coverage. To address this, Dey et al. (Dey et al., 2024) introduced a feedback-driven approach in which an autonomous agent refines search terms iteratively to mimic a journalist’s evolving research strategy. Such iterative refinements allow the system to capture elusive evidence more effectively than a single-pass search, particularly when no direct textual match exists for the claim.
2.2. Reasoning with evidence
@@
Data-driven fact-checking struggles with generalization, even when using external evidence. Prior supervised fact-checking models often overfit to specific patterns in training data and underperform on new domains. This was demonstrated by Tonglet et al. (Tonglet et al., 2025) who showed that state-of-the-art models (Papadopoulos et al., 2024b, a) degrade significantly on real-world datasets compared to curated benchmarks, likely due to domain mismatch. Moreover, even if external evidence is retrieved, many approaches simply score claim-evidence alignment in a single pass based on ranking, ignoring contradictory or ambiguous data. In contrast to the use of heuristics to rank evidence (Papadopoulos et al., 2024a), we propose clustering distinct narratives and applying judgment that accounts for support or conflict from each cluster.
@@
Figure 2. Overview of CRAVE: Given a claim (image + text), contained in a social media post, we retrieve evidence from reverse image and text-based searches. We then cluster the results into distinct narratives about the claim, selecting only visually relevant evidence (highlighted in red). Each narrative is compared to the claim in a prompt-based LLM to predict veracity and produce a judgment and explanation. Zoom in for details.
@@
LLM prompt-based veracity judgment has been explored more recently, incorporating both example demonstrations and fine-tuning on training sets. Qi et al. (Qi et al., 2024) predicts veracity by detecting inconsistencies between the caption, image, and visual entities using a fine-tuned Multimodal Large Language Model (MLLM) InstructBLIP(Dai et al., 2023), and between the caption and text evidence using a frozen LLM (Zheng et al., 2023). The approach was extended to multiple debating LLMs in (Lakara et al., 2025). These approaches degrade when applied to images and entities with non-Western contexts that the MLLMs have not encountered during training. Thus (Tonglet et al., 2025) uses predicted context and claim text as input, employing a frozen LLAMA 3(Grattafiori et al., 2024) with few-shot demonstrations to predict veracity. They also experiment with an LLM (He et al., 2021) fine-tuned on (predicted context, ground truth veracity) training pairs. However, the authors note that training on synthetic out of context data, which follow a different distribution than the real-world data, worsens performance compared to direct comparison without learning. One insight from such work is that minimal learning may be more robust for final classification if one can structure how the evidence is compared. In CRAVE, we explore this and focus on structured claim-evidence comparison by combining narrative clustering and LLM judgment. The concept of judgment based on different narratives is explored in (Khan et al., 2024), by having expert models argue for different answers. However, multi-turn debates pose the risk of one side dominating and unduly influencing the decision. In our approach, we simplify the process by using a single-step LLM judge that compares the generated narratives from the previous stage. This approach allows the judge to maintain greater control over the arguments considered, preventing any single viewpoint from influencing the outcome through prolonged interaction.
@@
In summary, current research primarily focuses on debunking or verifying claims using ranked retrieved evidence. However, for divisive content, where different media outlets present conflicting narratives, fact-checking systems often fall short by failing to accommodate multiple perspectives. Community notes on X (formerly Twitter) illustrate how different user-sourced evidence can be used to support opposing views, leading to debate and engagement rather than a definitive verdict(De et al., 2024). Arguably such an approach supports democratic discourse (Walter and Salovich, 2021), but has not been leveraged in automated fact-checking work. CRAVE addresses this gap by using cluster-based, multimodal retrieval that can group evidence into coherent narratives, enabling LLMs to more effectively reason over long or conflicting inputs.
3. Cluster-based Retrieval Augmented Verification with Explanation
@@
The objective of CRAVE (Fig. 2) is to predict and explain the veracity of a multimodal social media post $P$, which comprises an image $I_{P}$ and accompanying text $T_{P}$, by leveraging external evidence. Our approach retrieves relevant evidence, organizes it into coherent narratives, and delivers a final veracity judgment. The approach comprises four key stages: evidence retrieval, clustering and narrative extraction, agent-based evidence refinement, and explainable judgment.
3.1. Evidence retrieval
@@
A reverse image search is conducted using the input image $I_{P}$, retrieving a set of evidence pairs $\mathcal{X}_{I}=\{(I_{i},T_{i})\}_{i=1}^{n_{I}}$, where each pair $(I_{i},T_{i})$ comprises an image $I_{i}$ and its associated text $T_{i}$. Similar to (Abdelnabi et al., 2022; Dey et al., 2024; Tonglet et al., 2025), we use the Vision API(Google, 2025a) to perform reverse search. While (Abdelnabi et al., 2022) only uses the text entries from this reverse search, we use both returned evidence images and texts, similar to (Dey et al., 2024). This forms the image-based evidence set $\mathcal{X}_{I}$ that represents the image context.
@@
The input text $T_{P}$ is used to generate search queries with the assistance of an LLM which is applied to the Programmable search API (Google, 2025b). While the use of Programmable search has been seen in the past in (Abdelnabi et al., 2022; Dey et al., 2024; Tonglet et al., 2025), we take forward the idea of generating search terms and iterative searches introduced in (Dey et al., 2024). These queries retrieve an additional set of pairs of evidence $\mathcal{X}_{T}=\{(I_{j},T_{j})\}_{j=1}^{n_{T}}$, where each pair consists of an image $I_{j}$ and an associated text $T_{j}$. Our text-based search improves upon its initial retrievals through comparison of the evidence text with the claim text, to refine the search term for H subsequent iterations leveraging an LLM to directly discover inconsistencies. We determine the inconsistency in terms of extracted named entities related to 5W1H (Who, What, When, Where, Why, and How) (Urbani, 2020; Tonglet et al., 2024, 2025), with the search terms of the next iteration focused on entities not present in results from the current retrieval iteration. We try this for H=3 iterations, stopping before if we cannot find discrepancies in terms of 5W1H. This forms the text-based evidence set $\mathcal{X}_{T}$.
@@
3.2. Clustering and narrative extraction
@@
The retrieved evidence set $\mathcal{X}$ is clustered into $K=4$ distinct groups using the K-means algorithm. Each piece of evidence from $\mathcal{X}$ is first embedded into 384-dimensional dense vectors using Sentence Transformer (Reimers and Gurevych, 2019), applied to its text component $T$. Our insights indicate that incorporating multimodal representations, such as image embeddings, could lead to degradation in classification performance due to the frequent pairing of identical images with conflicting text narratives. Each cluster $\mathcal{C}_{k}$ is then defined as:
@@
The choice of $K$ was determined through experimentation on 300 samples from the Real-world fake news benchmark Verite, as later analyzed in Table 3. These coherent clusters enable retrieval of targeted evidence in Sec. 3.3 and also allows us to deal with noisy web data during judgment, detailed in Sec. 3.4.
3.3. Agent-based evidence refinement
@@
3.4. Explainable judgment
The refined clusters $\{\mathcal{C}_{k}^{\prime}\}_{k=1}^{K}$ and their associated narratives $\{N_{k}\}_{k=1}^{K}$ are evaluated by an LLM-based judge along with the original social media post $P$, to produce a veracity verdict and explanation. The verdict categories are:

### v2

@@ Abstract
I Introduction
Fact-checking has emerged as a critical task in the era of social media [1], where information spreads rapidly [2, 3] and misinformation can have far-reaching consequences [4, 5, 6]. Automated fact-checking systems have gained traction as scalable solutions, yet they often face challenges such as handling diverse evidence sources, integrating multimodal data, and presenting comprehensive narratives. Traditional approaches [7, 8] focus on verifying factual claims against a static repository of information [9, 10, 11, 12], which may overlook nuances in evidence and the broader social context in which a claim spreads [13].
Recent advances in Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) have opened new avenues for building robust fact-checking frameworks. RAG has been shown to be effective in reducing factual inconsistencies (”hallucinations”) during text generation, making it attractive for the fact-checking task [14, 15]; this is a crucial advantage, as such errors have been shown to significantly erode user trust in AI systems [16]. RAG models can access dynamic external knowledge, reason over complex evidence, and produce context-aware summaries. However, effectively leveraging these capabilities requires addressing several challenges: (1) how to retrieve relevant evidence efficiently, (2) how to cluster and organize potentially contradictory bodies of evidence into coherent narratives, and (3) how to generate holistic evaluations that incorporate these multiple perspectives.
Figure 1: CRAVE is a fact-checking method that analyses multimodal claims, expressed in social media posts, to determine their veracity. It retrieves multimodal evidence, clusters it into distinct narratives, and uses an LLM with 5W1H (who, what, when, where, why, and how) reasoning to produce an interpretable explanation and veracity verdict.
In this work, we propose CRAVE (Cluster-based Retrieval Augmented Verification with Explanation); a novel framework, Figure 1, that integrates retrieval-augmented LLMs with clustering techniques to identify, structure, and evaluate evidence for fact-checking social media posts. The framework is designed to process multi-modal inputs (text and images) and iteratively refine evidence through agent-based mechanisms. By clustering evidence into distinct narratives and identifying external support (or otherwise) for those narratives, CRAVE enables the assessment of contentious topics that present with support from contradictory and noisy news sources. The contributions of CRAVE are threefold:
@@ II Related work
Automated fact-checking systems deployed in practice commonly work by matching newly surfaced claims against a library of already fact-checked claims [17]. While efficient for well-known misinformation that periodically resurfaces, such systems struggle with novel or emerging misinformation. Research to date has mostly focused upon detecting intrinsic artifacts within a claim [18, 19] using synthetically generated datasets [20, 19, 18]. Such black-box detection methods often lack the kind of social and contextual explainability [1] needed for users to trust their verdicts [13], while also lagging behind developments in generative AI, limiting their effectiveness and on novel manipulation schemes.
Thus, while the detection of intrinsic artifacts within manipulated media—such as image edits, swaps [19], or deepfakes [18]—constitutes a valuable and active line of inquiry, we advocate for a paradigm shift towards external evidence-based verification. Rather than focusing on increasingly brittle and often ephemeral manipulation signatures within a given image, our approach prioritizes a comparative analysis against a corpus of external, verifiable sources. By leveraging robust image similarity measures, we can retrieve visually analogous content and, critically, analyze the surrounding context. This methodology allows for a more nuanced judgment: an image can be authenticated as “used in context,” flagged as “out of context” (a prevalent form of misinformation), or identified as a likely fabrication (e.g., a deepfake) due to its conspicuous absence from trusted external sources or its prior identification and debunking within them. This external, context-aware framework offers a more resilient and interpretable alternative to the current adversarial dynamic of detection and evasion.
@@
Evidence-driven methods and datasets have been introduced to enhance the explainability of fact-checking frameworks. However, the evidence used is often curated and sanitized [9, 11, 21]. As a result, methods such as [8, 7] achieve state-of-the-art results while bypassing the challenge of retrieving relevant evidence, which limits their applicability to real-world scenarios [1]. Although these approaches are evidence-based, they are primarily data-driven trained classifiers that struggle to generalize across novel claims.
II-A Retrieval of evidence
@@
Effective verification of multimodal misinformation depends on locating reliable evidence, particularly for images reused out of context [11, 7, 22]. As identified in [17] evidence retrieval is still done primarily by human fact-checkers through manually constructing search terms. Prior works have demonstrated that incorporating reverse image search to find an image’s provenance and text-based search for contextual information can similarly improve automated fact-checking performance [11, 12, 23, 24]. Abdelnabi et al. were among the first to show that combining reverse image search with text retrieval improves the scrutiny of multimedia claims. Since then, other pipelines have extended such concepts to incorporate more robust retrieval modules or specialized knowledge. In [12] the authors extend their sources through WIKIDATA [25] and Wikichat [26] to enrich their image context, but the authors also note the limited effectiveness in non-western contexts.
@@
In real-world scenarios, fact-checking is complicated by how social media users often combine verbose or biased commentary with images unrelated to mainstream coverage. To address this, Dey et al. [23] introduced a feedback-driven approach in which an autonomous agent refines search terms iteratively to mimic a journalist’s evolving research strategy. Such iterative refinements allow the system to capture elusive evidence more effectively than a single-pass search, particularly when no direct textual match exists for the claim.
II-B Reasoning with evidence
@@
Data-driven fact-checking struggles with generalization, even when using external evidence. Prior supervised fact-checking models often overfit to specific patterns in training data and underperform on new domains. This was demonstrated by Tonglet et al. [12] who showed that state-of-the-art models [8, 7] degrade significantly on real-world datasets compared to curated benchmarks, likely due to domain mismatch. Moreover, even if external evidence is retrieved, many approaches simply score claim-evidence alignment in a single pass based on ranking, ignoring contradictory or ambiguous data. In contrast to the use of heuristics to rank evidence [7], we propose clustering distinct narratives and applying judgment that accounts for support or conflict from each cluster.
@@
Figure 2: Overview of CRAVE: Given a claim (image + text), contained in a social media post, we retrieve evidence from reverse image and text-based searches. We then cluster the results into distinct narratives about the claim, selecting only visually relevant evidence (highlighted in red). Each narrative is compared to the claim in a prompt-based LLM to predict veracity and produce a judgment and explanation. In this example the false claim—that the IAF plane was shot down by Pakistan—was countered by two clusters: C1 showing it was a training crash (true context), while C3 included unrelated real incidents of downed planes but crucially also a fact-checking article debunking the claim. This shows that the system can identify true context (C1) and legitimate debunks (C3).
@@
LLM prompt-based veracity judgment has been explored more recently, incorporating both example demonstrations and fine-tuning on training sets. Qi et al. [24] predicts veracity by detecting inconsistencies between the caption, image, and visual entities using a fine-tuned Multimodal Large Language Model (MLLM) InstructBLIP[27], and between the caption and text evidence using a frozen LLM [28]. The approach was extended to multiple debating LLMs in [29]. These approaches degrade when applied to images and entities with non-Western contexts that the MLLMs have not encountered during training. Thus [12] uses predicted context and claim text as input, employing a frozen LLAMA 3 [30] with few-shot demonstrations to predict veracity. They also experiment with an LLM [31] fine-tuned on (predicted context, ground truth veracity) training pairs. However, the authors note that training on synthetic out-of-context data, which follow a different distribution than the real-world data, worsens performance compared to direct comparison without learning. One insight from such work is that minimal learning may be more robust for final classification if one can structure how the evidence is compared. In CRAVE, we explore this and focus on structured claim-evidence comparison by combining narrative clustering and LLM judgment. The concept of judgment based on different narratives is explored in [32], by having expert models argue for different answers. However, multi-turn debates pose the risk of one side dominating and unduly influencing the decision. In our approach, we simplify the process by using a single-step LLM judge that compares the generated narratives from the previous stage. This approach allows the judge to maintain greater control over the arguments considered, preventing any single viewpoint from influencing the outcome through prolonged interaction.
@@
In summary, current research primarily focuses on debunking or verifying claims using ranked retrieved evidence. However, for divisive content, where different media outlets present conflicting narratives, fact-checking systems often fall short by failing to accommodate multiple perspectives. Community notes on X (formerly Twitter) illustrate how different user-sourced evidence can be used to support opposing views, leading to debate and engagement rather than a definitive verdict [33]. Arguably such an approach supports democratic discourse [34], but has not been leveraged in automated fact-checking work. CRAVE addresses this gap by using cluster-based, multimodal retrieval that can group evidence into coherent narratives, enabling LLMs to more effectively reason over long or conflicting inputs.
III Cluster-based Retrieval Augmented Verification with Explanation
@@
The objective of CRAVE (Figure 2) is to predict and explain the veracity of a multimodal social media post $P$, which comprises an image $I_{P}$ and accompanying text $T_{P}$, by leveraging external evidence. Our approach retrieves relevant evidence, organizes it into coherent narratives, and delivers a final veracity judgment with an explanation. The approach comprises three key stages: evidence retrieval, clustering, and LLM reasoning.
III-A Retrieval
@@
A reverse image search is conducted using the input image $I_{P}$, retrieving a set of evidence pairs $\mathcal{X}_{I}=\{(I_{i},T_{i})\}_{i=1}^{n_{I}}$, where each pair $(I_{i},T_{i})$ comprises an image $I_{i}$ and its associated text $T_{i}$. Similar to [11, 23, 12], we use the Vision API [35] to perform reverse search. While [11] only uses the text entries from this reverse search, we use both returned evidence images and texts, similar to [23]. This forms the image-based evidence set $\mathcal{X}_{I}$ that represents the image context.
@@
The input text $T_{P}$ is used to generate search queries with the assistance of an LLM which is applied to the Programmable search API [36]. While the use of Programmable search has been seen in the past in [11, 23, 12], we take forward the idea of generating search terms and iterative searches introduced in [23]. These queries retrieve an additional set of pairs of evidence $\mathcal{X}_{T}=\{(I_{j},T_{j})\}_{j=1}^{n_{T}}$, where each pair consists of an image $I_{j}$ and an associated text $T_{j}$. Our text-based search improves upon its initial retrievals through comparison of the evidence text with the claim text, to refine the search term for H subsequent iterations leveraging an LLM to directly discover inconsistencies. We determine the inconsistency in terms of extracted named entities related to 5W1H (Who, What, When, Where, Why, and How) [37, 38, 12], with the search terms of the next iteration focused on entities not present in results from the current retrieval iteration. We try this for H=3 iterations, stopping before if we cannot find discrepancies in terms of 5W1H. This process augments the evidence set. This refinement identifies missing 5W1H information in a evidence and generates targeted new queries to find it, adding more relevant evidence. This forms the text-based evidence set $\mathcal{X}_{T}$.
@@
III-B Clustering: dealing with noisy evidence
Web-scraped evidence is often noisy and may not directly relate to the claim. This issue is exacerbated in older datasets, where images are frequently repurposed in different contexts across platforms. For verbose social media posts, retrieving relevant evidence relies on extracting factual search terms from claims. Additionally, if a claim’s images and text do not originate from published news articles, retrieval becomes more challenging since such content is rarely indexed by search engines. This challenge is further aggravated when the post is not in English or does not concern prominent Western contexts. Prior work highlights these limitations, such as for Spanish tweets [23] and image contextualization in East Africa and South Asia [12]. To handle noisy evidence, we employ two main strategies: text-based clustering and visual filtering.
III-B1 Text-based clustering
Texts with different narratives coupled with irrelevant texts—whether due to scraping failures or website restrictions—can introduce noise during decision-making. To address this, we use narrative clustering to group texts about related topics into similar clusters. This helps maintain distinct narratives while preventing conflicted narratives from being mixed.
Initial clustering and narrative extraction
@@ III-B1 Text-based clustering
The retrieved evidence set $\mathcal{X}$ is clustered into $K$ distinct groups using the K-means algorithm. Each piece of evidence from $\mathcal{X}$ is first embedded into 384-dimensional dense vectors using Sentence Transformer [39], applied to its text component $T$. Our insights indicate that incorporating multimodal representations, such as image embeddings, could lead to degradation in classification performance due to the frequent pairing of identical images with conflicting text narratives. Each cluster $\mathcal{C}_{k}$ is then defined as:
@@
The choice of $K=4$ was determined through experimentation on Real-world fake news benchmark Verite, 5pilsOOC, and DP, as analyzed in subsubsection V-C3, while the clustering method was determined through experimentation on random subsets of 300 samples each from these datasets, analyzed in subsubsection V-C4. These coherent clusters not only allow us to deal with noisy web data during judgment but also enable retrieval of targeted evidence detailed in subsubsection III-B1.
Agent-based cluster refinement
@@
III-B2 Visual filtering
Low visual similarity reliably indicates irrelevant evidence, thus we filter evidence using an image similarity threshold, following prior approaches [12, 23]. Our visual features are extracted from pre-trained networks that capture key visual aspects: faces, scenes, and visual semantics. Faces are detected using [40], with the most prominent face embedded via [41], yielding $\mathcal{V}^{face}\in\mathbb{R}^{512}$. Scene information is extracted via a model trained on 365 place categories [42], producing $\mathcal{V}^{place}\in\mathbb{R}^{2048}$. Visual semantics are captured using [43], resulting in $\mathcal{V}^{sem}\in\mathbb{R}^{1000}$. The final visual representation of an image is:
$\mathcal{V}=[\mathcal{V}^{face},\mathcal{V}^{place},\mathcal{V}^{sem}].$
The cosine similarity between $\mathcal{V}^{I_{P}}$ and $\mathcal{V}^{I_{k}}$ defines the final similarity score of an evidence image $I_{k}$in $\mathcal{C}_{k}^{\prime}$ with the claim image $I_{P}$. We compute cosine similarity for each vector type (face, place, semantics) separately and then combine with equal weighting, avoiding dimensionality issues. Following [12] we leverage the large training set of NewsCLIPpings to determine our similarity threshold of 0.9.
Since reverse image searches are designed to retrieve high-quality, contextually relevant images, evidence with low visual similarity—typically originating from text-based searches—can be safely disregarded, as it is unlikely to offer meaningful context for verification. However, visually similar yet textually different evidence may indicate out-of-context usage rather than irrelevance. Thus by using visual similarity for filtering and clustering noisy textual evidence, we ensure that relevant evidence, even if with a different text context, is retained and presented in a structured manner to the LLM for reasoning.
LLM-based veracity reasoning Inputs: Claim text: $T_{P}$ Clustered evidence: $\left\{\mathcal{C}_{k}^{\prime}\right\}_{k=1}^{K}$, with each $\mathcal{C}_{k}^{\prime}$ having: Direct text search evidence: $\mathcal{X}_{T_{k}}$ Reverse image search evidence: $\mathcal{X}_{I_{k}}$ Narrative assessment: 1. For each evidence cluster $\mathcal{C}_{k}^{\prime}$, the LLM compares the claim $T_{P}$ with the text component of $\mathcal{X}_{T_{k}}$ and $\mathcal{X}_{I_{k}}$ in terms of Location, Named person, Date, Main topic and Common objects 2. Based on alignment, with $\mathcal{X}_{T_{k}}$ and $\mathcal{X}_{I_{k}}$, each Narrative $\mathcal{C}_{k}^{\prime}$ is labeled as: Supports – entities align and no conflict Conflicts – misalignment in entities Irrelevant – lacks named entities Reasoning with narratives: 3. LLM applies a decision strategy prioritizing $\mathcal{X}_{I_{k}}$: If any Narrative with valid $\mathcal{X}_{I_{k}}$ supports the claim
→ label as True
If Narratives with $\mathcal{X}_{I_{k}}$ conflict or are irrelevant
→Check Narratives with $\mathcal{X}_{T_{k}}$
If any Narratives with $\mathcal{X}_{T_{k}}$ supports
→ label as True
If all Narratives with $\mathcal{X}_{T_{k}}$ conflict or are irrelevant
→ label as Misleading
4. The LLM generates a textual justification grounded in the above alignment assessments.
Figure 3: Reasoning protocol used by the LLM (GPT 4o) to assess the veracity of a post $(T_{P},I_{P})$ using clustered direct ($\mathcal{X}_{T_{k}}$) and reverse image search ($\mathcal{X}_{I_{k}}$) evidence. Narrative assessment includes comparison and labeling, while reasoning prioritizes reverse image evidence to determine a final verdict.
III-C Reasoning: prompt-based comparison and explanation
The refined clusters $\{\mathcal{C}_{k}^{\prime}\}_{k=1}^{K}$ and their associated narratives $\{N_{k}\}_{k=1}^{K}$ are evaluated by an LLM-based judge along with the original social media post $P$, to produce a veracity verdict and explanation.
The verdict categories are: